OpenStudy (anonymous):

If ax*x + bx +c = 0, then what is x?

OpenStudy (anonymous):

(a)x*x + (b)x + (c) = 0 d*d = b*b - 4ac (-b +/- d)/2a=x

OpenStudy (anonymous):

k thanks have lots of hw

OpenStudy (anonymous):

No problem. And I know how yo u feel. I had to finish four pages of ! algebra 2hw last night:/

OpenStudy (anonymous):

k thanks have lots of hw

OpenStudy (anonymous):

oh.... thats a lot!